# On the Pitfalls of Batch Normalization for End-to-End Video Learning: A Study on Surgical Workflow Analysis

### v4

@@ 3.2 BatchNorm’s Problems with Sequential Data
1) Small-batch effect (Sec. 4.2, Fig. 33): It is well-known that BatchNorm performs poorly with small batches (Wu and He, 2018; Wu and Johnson, 2021). Batch statistics, used to normalize features during training, only poorly estimate global statistics used during inference and thus induce a discrepancy between train and test behavior. However, when batches contain sequential video frames, even large batches lead to similar issues as the highly correlated samples also do not approximate global statistics well. Similar observations have been made in different contexts (Ba et al., 2016; Wu and He, 2018).
2) Cheating/leakage (Sec.4.3, Fig. 67): Previous work has discussed BatchNorm’s ability to leak information from other samples in the batch and its impact on performance in object detection or contrastive learning (Wu and Johnson, 2021). In online video tasks, a similar phenomenon can be fatal. BatchNorm can leak information from future frames, which allows models to “cheat” certain objectives and prevents learning useful features. This “cheating” phenomenon is most obvious in anticipation tasks. In instrument anticipation, the occurrence of an instrument influences the batch statistics of channels reacting to its appearance. During training, BN models may pick up on these fluctuations in batch statistics even at earlier time points within the same batch and use them to anticipate the instrument’s later occurrence. Thus, the model only has to learn to recognize the instrument to solve the anticipation objective during training, but will fail during inference when batch statistics are not available.
@@ 3.3 Alternative Backbones without BatchNorm
We enable end-to-end learning with single-sequence batches by replacing ResNet50 with existing BN-free CNNs. We use a ResNet50 with GroupNorm (GN) (Wu and He, 2018) for fair comparison as well as the more recent ConvNeXt-T (Liu et al., 2022b), which uses ViT-style LayerNorm (Yao et al., 2021a) and is of comparable size to ResNet50. We also investigate FrozenBN (Liu et al., 2022a; Wu and Johnson, 2021), where batch statistics are replaced by global statistics from the pretraining stage. All models are pretrained on ImageNet-1k (Krizhevsky et al., 2012).
@@ 3.4 Limitations of Previous Strategies
Multi-stage: Multi-stage training is the most popular approach and effectively avoids BN’s issues by training backbones without temporal context. Yet, this has several disadvantages. Firstly, it increases the number of hyperparameters since learning rate, number of epochs etc. have to be tuned for each training stage. Secondly, when backbone and temporal model are trained on the same data, overfitting on the image task can prevent the temporal model from learning useful features. For example, we found early stopping of the ResNet pretraining crucial to reproduce TeCNO’s (Czempiel et al., 2020) results. Lastly and most importantly, pretraining on single frames changes the dynamics of the task and cannot guarantee features that are suitable for the target task. A certain visual feature might only be relevant to the task in combination with a previous event. If the backbone does not learn to capture this feature during pretraining, this information is lost and cannot be recovered in the subsequent temporal-modeling stage.
End-to-end with multi-sequence batches: Sampling multiple sequences per batch (e.g. SV-RCNet (Jin et al., 2017)) reduces the discrepancy between batch and global statistics. However, this comes at the cost of shorter training sequences. It also restricts possible training strategies like carrying hidden states of LSTMs across batches and can thus impair performance.
@@ 4 Experiments
Fig. 2: Phase Recognition (Cholec80): BN-free, end-to-end CNN-LSTMs outperform 2-stage LSTMs and TCNs. The widely-used ResNet50 with BN fails. Split: 40/8/32 from TeCNO (Czempiel et al., 2020).
Refer to caption (a) Length vs diversity Refer to caption (b) CHT strategy improves BN-free models Refer to caption (c) Partial freezing
Fig. 2: Phase Recognition (Cholec80): BN-free, end-to-end CNN-LSTMs outperform 2-stage LSTMs and TCNs. The widely-used ResNet50 with BN fails. Split: 40/8/32 from TeCNO (Czempiel et al., 2020).
Fig. 3: Phase Recognition (Cholec80): (a) While longer sequences improve BN-free models, BN in ResNet50 backbones requires careful selection of the number of sequences per batch before performance drops. (b) Carrying hidden state through entire videos at test time improves models. By also doing this during training, train-test discrepancy is removed and models improve further. In BN models, implementing this is not straightforward. (c) BN models improve by freezing backbone blocks and thus increasing training sequences while maintaining batch diversity. Yet, freezing also improves BN-free models. Explicit scores with standard deviations and additional metrics can be found in the appendix (Tables 7 & 8).
@@ 4.1 Data, Tasks & Implementation Details
Cholec80: For our main analysis, we train phase recognition and anticipation models on the widely-studied Cholec80 dataset (Twinanda et al., 2016), which consists of 80 recorded gallbladder removals (cholecystectomies). Videos range from 12 to 100 min (mean ca. 38 min), processed at 1fps with frame-wise labels for 7 instruments and 7 surgical phases.
Surgical phase recognition is an dense online temporal action segmentation task (Twinanda et al., 2016). We use the mean video-wise accuracy for evaluation (F1 in Appendix Table 8).
Instrument anticipation is an online frame-wise regression task. At each time point, the objective is to predict the remaining time until occurrence of an instrument within a horizon of 5 minutes. Outside the horizon, a constant should be predicted (see Fig. 1). We follow the recently proposed task formulation (Rivoir et al., 2020). We use the weighted MAE (wMAE) as evaluation metric, where errors inside and outside the horizon are weighted equally to counteract imbalances.
For both tasks, we use 40 videos for training, 8 for validation and 32 for testing, following TeCNO (Czempiel et al., 2020). Epochs with the best validation score (acc./wMAE) are selected for testing. For SOTA comparison in phase recognition, we retrain models on a 32/8/40 split. This allows us to stay comparable to the widely used 40/40 split (Twinanda et al., 2016; Gao et al., 2021) while still having a validation set for model selection. All runs are repeated 3 times. More training and evaluation details can be found in E & F.
GTEA & 50Salads: To reproduce our claims on natural videos, we retrain phase recognition models for online action segmentation on GTEA (Fathi et al., 2011) and 50Salads (Stein and McKenna, 2013).
AutoLaparo & CATARACTS: For state-of-the-art comparisons, we retrain phase recognition models on two additional surgical datasets. AutoLaparo (Wang et al., 2022) contains 21 long videos (10/4/7 split) ranging from 27 to 112 min (mean ca. 66 min) with 7 coarse phases annotated. CATARACTS (Zisimopoulos et al., 2018) contains 50 short videos ranging from 6 to 40 min (mean ca. 11 min) with 19 fine-grained steps annotated. To be able to compare to the 25/25 split used in previous work (Chen et al., 2022), we use a 20/5/25 split. On both datasets, we report the mean video-based accuracy as well as macro-precision, -recall and -jaccard, which are computed over all test frames. See E & F for details.
@@ 4.2 Surgical Phase Recognition
Hypothesis 1: End-to-end outperforms 2-stage training, but not with BatchNorm. We argue that end-to-end learning is preferable over 2-stage approaches but fails when the backbone contains BatchNorm. To demonstrate this, we train CNN-LSTM models with three different backbones (Resnet50-BN, ResNet50-GN and ConvNeXt-T), single-sequence batches of 64 frames ($1\times 64$) and the proposed carry-hidden training (CHT). For comparison, we train each backbone in a 2-stage process with an LSTM or causal MS-TCN, following TeCNO’s training procedure (Czempiel et al., 2020).
Fig. 3 shows that end-to-end models outperform 2-stage approaches when using a backbone without BN. Only the typically used BN-based ResNet50 fails completely. While BatchNorm is known to fail with small batch sizes (Wu and He, 2018; Wu and Johnson, 2021), even large batch sizes lead to similar issues here due to the strong correlation of subsequent frames. In other words, a batch with a single 64-frame sequence behaves similar to a batch size of 1. The next hypothesis investigates this in more detail.
Hypothesis 2: BatchNorm causes trade-off between sequence length and batch diversity. We investigate the effect of different batch configurations and follow SV-RCNet’s (Jin et al., 2017) strategy for end-to-end CNN-LSTMs: Multiple shorter sequences are sampled per batch, hidden states are reset after each training sequence and the LSTM uses a sliding window for evaluation (SWE) to ensure consistent sequence lengths at train and test time. We test this strategy using 1-8 sequences in batches of 64 frames in total.
In Fig. 2(a), we observe that the performance of all BN-free models improves with fewer but longer sequences per batch since it increases the temporal context. In BN models, however, fewer sequences poorly approximate global statistics, leading to a discrepancy in feature normalization at train and test time. Thus, we observe drops in performance despite the constant total batch size, similar to the frame-rate trade-off in video-clip classification (Wu and He, 2018). The BN backbone outperforms its FrozenBN and GroupNorm counterparts for multiple short sequences but stagnates and then collapses with fewer, longer ones. Due to this trade-off between sequence length and batch diversity, BN models eventually cannot reach the performance of the fully-sequential, BN-free models.
@@
Hypothesis 3: BatchNorm restricts LSTM usage for online tasks. Note that by adopting SV-RCNet’s hidden-state handling, none of the methods outperform the 2-stage SOTA (Czempiel et al., 2020). This is likely due to the short temporal context seen by the LSTM. Simply carrying the hidden state across the entire video during inference (CHE) already provides a clear performance boost without changing the training strategy (Fig. 3, SWE vs. CHE). Yet, this induces another train-test discrepancy since the LSTM accumulates hidden states over much longer time periods at test time compared to training time, which could cause unpredictable behavior.
@@
Fig. 5: Online action segmentation: Results on surgical data (left) can be reproduced on natural-video datasets (right).
Hypothesis 4: BatchNorm induces similar issues on natural-video datasets. Fig. 5 confirms our findings on small-scale datasets from the natural-video domain (GTEA and 50Salads) by retraining the same models used for phase recognition on online action segmentation. We find that our main claims can be reproduced, namely that (1) BN-based end-to-end models fail with singöle-sequence batches, (2) end-to-end approaches without BN outperform 2-stage models and (3) the CHT strategy is effective using GN or ConvNeXt but not with FrozenBN. Our proposed strategies aim at increasing temporal context, and thus seem mostly applicable to tasks where this is relevant (e.g. long videos and workflow-based tasks such as cooking or other complex activities). Its impact on other video tasks remains open.
Hypothesis 5: Other BatchNorm variants are not suitable. In Table 2, we perform a more extensive comparison of batch-independent and batch-dependent normalization layers. We find that other batch-independent methods (InstanceNorm, LayerNorm) benefit from our context-enlarging strategies. However, even BatchNorm variants which attempt to alleviate issues with temporal dependence or small batches exhibit similar failure cases as BatchNorm. Temporal dependence can be avoided entirely by normalizing each time step individually. Yet, with single-sequence batches this still fails due to normalization over single frames. DNR (Cai et al., 2021) proposes temporally dependent estimation of BN’s affine parameters to adjust to properties of batches in video settings. Yet, batch stats are still computed like in standard BN, so the train/test discrepancy with single-sequence batches remains and model collapse with 1x64 batches can clearly be observed. CBN (Yao et al., 2021b) reduces batch dependence by accumulating stats from multiple batches. Yet, for our proposed CHT strategy, CBN is not suitable due to strong inter-batch correlations of subsequent batches. Further, we found the transition to accumulated stats after its burn-in phase to cause instability even in CHE settings, despite finetuning. Expectedly, this issue was stronger in the single-sequence case (1x64, CHE).
@@
Accuracy (mean±$\pm$±std. over runs) on 40/8/32 split from TeCNO (Czempiel et al., 2020).
@@
InstanceNorm (Ulyanov et al., 2016) $87.3(\pm 2.5)$ $88.5(\pm 0.5)$ $\bf 90.2(\pm 0.2)$
LayerNorm (Ba et al., 2016) $87.6(\pm 1.6)$ $89.1(\pm 1.3)$ $\bf 90.4(\pm 2.8)$
GroupNorm (Wu and He, 2018) $87.9(\pm 0.8)$ $89.3(\pm 0.4)$ $\bf 90.5(\pm 0.9)$
@@
DNR (Cai et al., 2021) $88.9(\pm 0.7)$ $68.0(\pm 6.5)$ $61.6(\pm 2.7)$
CBN (Yao et al., 2021b) w/ window size 1 (equiv. to BN) $89.5(\pm 0.7)$
@@
Through freezing, 4x64-BN finally outperforms the 2-stage SOTA but collapses again with longer sequences while GN models improve. Even FrozenBN matches other BN-free models. Consistent with prior beliefs (He et al., 2016), lack of normalization might be less problematic in shallow models.
Table 3: State of the art in surgical phase recognition on Cholec80 (Twinanda et al., 2016) with a 40/40 split. Our models are trained on a 32/8/40 split to allow for model selection. We find that simple, end-to-end CNN-LSTMs match or outperform multi-stage methods with attention-based temporal models (Gao et al., 2021; Jin et al., 2021; Chen et al., 2022), complex augmentation pipelines (Yi et al., 2022) or Transformer backbones (Chen et al., 2022; He et al., 2022).
@@
SV-RCNet (Jin et al., 2017) (E2E, 100x10, SWE) ResNet-BN LSTM 2 $85.3(\pm 7.3)$ $85.0$ $65.5$
TeCNO (Czempiel et al., 2020) ResNet-BN TCN 2 $88.6(\pm 7.8)$ $75.1(\pm 6.9)$ $88.1(\pm 0.3)^{\dagger}$ $69.4(\pm 0.5)^{\dagger}$ $85.4(\pm 1.0)^{\dagger}$
TMRNet (Jin et al., 2021) ResNet-BN LSTM, Attn. 2 $89.2(\pm 9.4)$ $78.9(\pm 5.8)$
Trans-SVNet (Gao et al., 2021) ResNet-BN TCN, Transf. 3 $90.1(\pm 7.1)$ $79.3(\pm 6.6)$ $89.0(\pm 1.1)^{\dagger}$ $71.5(\pm 2.4)^{\dagger}$ $85.4(\pm 1.4)^{\dagger}$
Not-E2E (Yi et al., 2022) ResNet-BN TCN, GRU 3 $92.0(\pm 5.3)$ $77.1(\pm 11.5)$
TMRNet (Jin et al., 2021) ResNeST LSTM, Attn. 2 $90.1(\pm 7.6)$ $79.1(\pm 5.7)$
PATG (Kadkhodamohammadi et al., 2022) SE-ResNet GNN 2 $91.4$
Zhang et al. (2022) C3D LSTM 2 $85.9(\pm 7.9)$
He et al. (2022) I3D GRU 2 $88.3(\pm 1.0)$
TeSTra (Zhao and Krähenbühl, 2022) Transf. Transf. 2 $90.1$ $71.6$
He et al. (2022) Video-Transf. GRU 2 $90.9(\pm 0.01)$
Dual Pyramid (Chen et al., 2022) Transf. Transf. 2 $91.4$ $75.4$
@@
LoViT (Liu et al., 2023a) Transf. Transf. 2 $\it 92.4(\pm 6.3)$ $\it 81.2(\pm 9.1)$ $\it 91.5$ $\it 74.2$
SKiT (Liu et al., 2023b) Transf. Transf. 2 $\it 93.4(\pm 5.2)$ $\it 82.6$ $\it 92.5$ $\it 76.7$
@@
Table 4: State of the art in surgical phase recognition on AutoLaparo (Wang et al., 2022) with 10/4/7 split. We report mean±$\pm$±std. over runs.
@@
SV-RCNet (Jin et al., 2017) $75.6$ $64.0$ $59.7$ $47.2$
TMRNet (Jin et al., 2021) $78.2$ $66.0$ $61.5$ $49.6$
TeCNO (Czempiel et al., 2020) $77.3$ $66.9$ $64.6$ $50.7$
Trans-SVNet (Gao et al., 2021) $78.3$ $64.2$ $62.1$ $50.7$
LoViT (Liu et al., 2023a) $81.4$ $\bf 85.1$ $65.9$ $56.0$
SKiT (Liu et al., 2023b) $82.9$ $81.8$ $70.1$ $59.9$
@@
Table 5: State of the art in surgical step recognition on CATARACTS (Zisimopoulos et al., 2018) with a 25/25 split. Our models are trained on a 20/5/25 split to allow for model selection. State-of-the-art scores are printed bold and we underline scores which outperform all methods except Chen et al. (2022).
@@
3D-CNN (Funke et al., 2019) $80.1$ $66.2$ $55.7$ $45.9$
SV-RCNet (Jin et al., 2017) $81.3$ $66.0$ $57.0$ $47.2$
TeCNO (Czempiel et al., 2020) $79.0$ $62.6$ $56.9$ $45.1$
Trans-SVNet (Gao et al., 2021) $77.8$ $61.3$ $55.0$ $43.8$
Dual Pyramid (Chen et al., 2022) $\bf 84.2$ $69.3$ $\bf 66.4$ $\bf 53.7$
@@
Cholec80: To be comparable to previous work (Gao et al., 2021; Jin et al., 2017, 2021), we retrain our models on a 32/8/40 split. As evaluation metrics, we report the mean video-wise accuracy and the mean video-wise balanced accuracy (i.e. mean video-wise macro recall) as well as variants of accuracy and Jaccard scores with relaxed boundaries, i.e. where errors at phase transitions are ignored under certain conditions. Relaxed metrics were proposed in the M2CAI 2016 Challenge (Stauder et al., 2016; Twinanda et al., 2016). Table 3 shows our results.
Through partial freezing on 256-frame sequences, all our models with BN-free backbones (FrozenBN, GN and ConvNeXt) outperform the previous state of the art, including complex multi-stage methods with attention-based temporal models (Gao et al., 2021; Jin et al., 2021; Chen et al., 2022), complex augmentation pipelines (Yi et al., 2022) or Transformer backbones (Chen et al., 2022; He et al., 2022). Models trained completely end to end (E2E) on shorter 64-frame sequences perform slightly worse. However, the ConvNeXt variant still achieves SOTA performance and the GN variant beats or matches other ResNet-based methods. Note that Not-E2E (Yi et al., 2022) achieves slightly higher relaxed accuracy but much lower relaxed Jaccard. The completely end-to-end FrozenBN model underperforms, as already observed in Hypothesis 3.

### v5

@@ 3.2 BatchNorm’s Problems with Sequential Data
1) Small-batch effect (Sec. 4.2, Fig. 33): It is well-known that BatchNorm performs poorly with small batches [76, 77]. Batch statistics, used to normalize features during training, only poorly estimate global statistics used during inference and thus induce a discrepancy between train and test behavior. However, when batches contain sequential video frames, even large batches lead to similar issues as the highly correlated samples also do not approximate global statistics well. Similar observations have been made in different contexts [4, 76].
2) Cheating/leakage (Sec.4.3, Fig. 67): Previous work has discussed BatchNorm’s ability to leak information from other samples in the batch and its impact on performance in object detection or contrastive learning [77]. In online video tasks, a similar phenomenon can be fatal. BatchNorm can leak information from future frames, which allows models to “cheat” certain objectives and prevents learning useful features. This “cheating” phenomenon is most obvious in anticipation tasks. In instrument anticipation, the occurrence of an instrument influences the batch statistics of channels reacting to its appearance. During training, BN models may pick up on these fluctuations in batch statistics even at earlier time points within the same batch and use them to anticipate the instrument’s later occurrence. Thus, the model only has to learn to recognize the instrument to solve the anticipation objective during training, but will fail during inference when batch statistics are not available.
@@ 3.3 Alternative Backbones without BatchNorm
We enable end-to-end learning with single-sequence batches by replacing ResNet50 with existing BN-free CNNs. We use a ResNet50 with GroupNorm (GN) [76] for fair comparison as well as the more recent ConvNeXt-T [50], which uses ViT-style LayerNorm [81] and is of comparable size to ResNet50. We also investigate FrozenBN [47, 77], where batch statistics are replaced by global statistics from the pretraining stage. All models are pretrained on ImageNet-1k [43].
@@ 3.4 Limitations of Previous Strategies
Multi-stage: Multi-stage training is the most popular approach and effectively avoids BN’s issues by training backbones without temporal context. Yet, this has several disadvantages. Firstly, it increases the number of hyperparameters since learning rate, number of epochs etc. have to be tuned for each training stage. Secondly, when backbone and temporal model are trained on the same data, overfitting on the image task can prevent the temporal model from learning useful features. For example, we found early stopping of the ResNet pretraining crucial to reproduce TeCNO’s [13] results. Lastly and most importantly, pretraining on single frames changes the dynamics of the task and cannot guarantee features that are suitable for the target task. A certain visual feature might only be relevant to the task in combination with a previous event. If the backbone does not learn to capture this feature during pretraining, this information is lost and cannot be recovered in the subsequent temporal-modeling stage.
End-to-end with multi-sequence batches: Sampling multiple sequences per batch (e.g. SV-RCNet [37]) reduces the discrepancy between batch and global statistics. However, this comes at the cost of shorter training sequences. It also restricts possible training strategies like carrying hidden states of LSTMs across batches and can thus impair performance.
@@ 4 Experiments
Fig. 2: Phase Recognition (Cholec80): BN-free, end-to-end CNN-LSTMs outperform 2-stage LSTMs and TCNs. The widely-used ResNet50 with BN fails. Split: 40/8/32 from TeCNO [13].
(a) Length vs diversity
(b) CHT strategy improves BN-free models
(c) Partial freezing
Fig. 3: Phase Recognition (Cholec80) with end-to-end CNN-LSTMs: (a) While longer sequences improve BN-free models, BN in ResNet50 backbones requires careful selection of the number of sequences per batch before performance drops. (b) Carrying hidden state through entire videos at test time improves models. By also doing this during training, train-test discrepancy is removed and models improve further. In BN models, implementing this is not straightforward. (c) BN models improve by freezing backbone blocks and thus increasing training sequences while maintaining batch diversity. Yet, freezing also improves BN-free models. Explicit scores with standard deviations and additional metrics can be found in the appendix (Tables 7 & 8).
@@ 4.1 Data, Tasks & Implementation Details
Cholec80: For our main analysis, we train phase recognition and anticipation models on the widely-studied Cholec80 dataset [70], which consists of 80 recorded gallbladder removals (cholecystectomies). Videos range from 12 to 100 min (mean ca. 38 min), processed at 1fps with frame-wise labels for 7 instruments and 7 surgical phases.
Surgical phase recognition is an dense online temporal action segmentation task [70]. We use the mean video-wise accuracy for evaluation (F1 in Appendix Table 8).
Instrument anticipation is an online frame-wise regression task. At each time point, the objective is to predict the remaining time until occurrence of an instrument within a horizon of 5 minutes. Outside the horizon, a constant should be predicted (see Fig. 1). We follow the recently proposed task formulation [58]. We use the weighted MAE (wMAE) as evaluation metric, where errors inside and outside the horizon are weighted equally to counteract imbalances.
For both tasks, we use 40 videos for training, 8 for validation and 32 for testing, following TeCNO [13]. Epochs with the best validation score (acc./wMAE) are selected for testing. For SOTA comparison in phase recognition, we retrain models on a 32/8/40 split. This allows us to stay comparable to the widely used 40/40 split [70, 24] while still having a validation set for model selection. All runs are repeated 3 times. More training and evaluation details can be found in E & F.
GTEA & 50Salads: To reproduce our claims on natural videos, we retrain phase recognition models for online action segmentation on GTEA [20] and 50Salads [67].
AutoLaparo & CATARACTS: For state-of-the-art comparisons, we retrain phase recognition models on two additional surgical datasets. AutoLaparo [75] contains 21 long videos (10/4/7 split) ranging from 27 to 112 min (mean ca. 66 min) with 7 coarse phases annotated. CATARACTS [91] contains 50 short videos ranging from 6 to 40 min (mean ca. 11 min) with 19 fine-grained steps annotated. To be able to compare to the 25/25 split used in previous work [12], we use a 20/5/25 split. On both datasets, we report the mean video-based accuracy as well as macro-precision, -recall and -jaccard, which are computed over all test frames. See E & F for details.
@@ 4.2 Surgical Phase Recognition
Hypothesis 1: End-to-end outperforms 2-stage training, but not with BatchNorm. We argue that end-to-end learning is preferable over 2-stage approaches but fails when the backbone contains BatchNorm. To demonstrate this, we train CNN-LSTM models with three different backbones (Resnet50-BN, ResNet50-GN and ConvNeXt-T), single-sequence batches of 64 frames ($1\times 64$) and the proposed carry-hidden training (CHT). For comparison, we train each backbone in a 2-stage process with an LSTM or causal MS-TCN, following TeCNO’s training procedure [13].
Fig. 3 shows that end-to-end models outperform 2-stage approaches when using a backbone without BN. Only the typically used BN-based ResNet50 fails completely. While BatchNorm is known to fail with small batch sizes [76, 77], even large batch sizes lead to similar issues here due to the strong correlation of subsequent frames. In other words, a batch with a single 64-frame sequence behaves similar to a batch size of 1. The next hypothesis investigates this in more detail.
Hypothesis 2: BatchNorm causes trade-off between sequence length and batch diversity. We investigate the effect of different batch configurations and follow SV-RCNet’s [37] strategy for end-to-end CNN-LSTMs: Multiple shorter sequences are sampled per batch, hidden states are reset after each training sequence and the LSTM uses a sliding window for evaluation (SWE) to ensure consistent sequence lengths at train and test time. We test this strategy using 1-8 sequences in batches of 64 frames in total.
In Fig. 3(a), we observe that the performance of all BN-free models improves with fewer but longer sequences per batch since it increases the temporal context. In BN models, however, fewer sequences poorly approximate global statistics, leading to a discrepancy in feature normalization at train and test time. Thus, we observe drops in performance despite the constant total batch size, similar to the frame-rate trade-off in video-clip classification [76]. The BN backbone outperforms its FrozenBN and GroupNorm counterparts for multiple short sequences but stagnates and then collapses with fewer, longer ones. Due to this trade-off between sequence length and batch diversity, BN models eventually cannot reach the performance of the fully-sequential, BN-free models.
@@
Hypothesis 3: BatchNorm restricts LSTM usage for online tasks. Note that by adopting SV-RCNet’s hidden-state handling, none of the methods outperform the 2-stage SOTA [13]. This is likely due to the short temporal context seen by the LSTM. Simply carrying the hidden state across the entire video during inference (CHE) already provides a clear performance boost without changing the training strategy (Fig. 3, SWE vs. CHE). Yet, this induces another train-test discrepancy since the LSTM accumulates hidden states over much longer time periods at test time compared to training time, which could cause unpredictable behavior.
@@
Fig. 5: Online action segmentation with end-to-end CNN-LSTMs: Results on surgical data (left) can be reproduced on natural-video datasets (right).
Hypothesis 4: BatchNorm induces similar issues on natural-video datasets. Fig. 5 confirms our findings on small-scale datasets from the natural-video domain (GTEA and 50Salads) by retraining the same models used for phase recognition on online action segmentation. We find that our main claims can be reproduced, namely that (1) BN-based end-to-end models fail with single-sequence batches, (2) end-to-end approaches without BN outperform 2-stage models and (3) the CHT strategy is effective using GN or ConvNeXt but not with FrozenBN. Our proposed strategies aim at increasing temporal context, and thus seem mostly applicable to tasks where this is relevant (e.g. long videos and workflow-based tasks such as cooking or other complex activities). Its impact on other video tasks remains open.
Hypothesis 5: Other BatchNorm variants are not suitable. In Table 2, we perform a more extensive comparison of batch-independent and batch-dependent normalization layers. We find that other batch-independent methods (InstanceNorm, LayerNorm) benefit from our context-enlarging strategies. However, even BatchNorm variants which attempt to alleviate issues with temporal dependence or small batches exhibit similar failure cases as BatchNorm. Temporal dependence can be avoided entirely by normalizing each time step individually. Yet, with single-sequence batches this still fails due to normalization over single frames. DNR [10] proposes temporally dependent estimation of BN’s affine parameters to adjust to properties of batches in video settings. Yet, batch stats are still computed like in standard BN, so the train/test discrepancy with single-sequence batches remains and model collapse with 1x64 batches can clearly be observed. CBN [82] reduces batch dependence by accumulating stats from multiple batches. Yet, for our proposed CHT strategy, CBN is not suitable due to strong inter-batch correlations of subsequent batches. Further, we found the transition to accumulated stats after its burn-in phase to cause instability even in CHE settings, despite finetuning. Expectedly, this issue was stronger in the single-sequence case (1x64, CHE).
@@
Accuracy (mean±$\pm$±std. over runs) on 40/8/32 split from TeCNO [13].
@@
InstanceNorm [72] $87.3(\pm 2.5)$ $88.5(\pm 0.5)$ $\bf 90.2(\pm 0.2)$
LayerNorm [4] $87.6(\pm 1.6)$ $89.1(\pm 1.3)$ $\bf 90.4(\pm 2.8)$
GroupNorm [76] $87.9(\pm 0.8)$ $89.3(\pm 0.4)$ $\bf 90.5(\pm 0.9)$
@@
DNR [10] $88.9(\pm 0.7)$ $68.0(\pm 6.5)$ $61.6(\pm 2.7)$
CBN [82] w/ window size 1 (equiv. to BN) $89.5(\pm 0.7)$
@@
Through freezing, 4x64-BN finally outperforms the 2-stage SOTA but collapses again with longer sequences while GN models improve. Even FrozenBN matches other BN-free models. Consistent with prior beliefs [27], lack of normalization might be less problematic in shallow models.
Table 3: State of the art in surgical phase recognition on Cholec80 [70] with a 40/40 split. Our models are trained on a 32/8/40 split to allow for model selection. We find that simple, end-to-end CNN-LSTMs match or outperform multi-stage methods with attention-based temporal models [24, 39, 12], complex augmentation pipelines [85] or Transformer backbones [12, 28].
@@
SV-RCNet [37] (E2E, 100x10, SWE) ResNet-BN LSTM 2 $85.3(\pm 7.3)$ $85.0$ $65.5$
TeCNO [13] ResNet-BN TCN 2 $88.6(\pm 7.8)$ $75.1(\pm 6.9)$ $88.1(\pm 0.3)^{\dagger}$ $69.4(\pm 0.5)^{\dagger}$ $85.4(\pm 1.0)^{\dagger}$
TMRNet [39] ResNet-BN LSTM, Attn. 2 $89.2(\pm 9.4)$ $78.9(\pm 5.8)$
Trans-SVNet [24] ResNet-BN TCN, Transf. 3 $90.1(\pm 7.1)$ $79.3(\pm 6.6)$ $89.0(\pm 1.1)^{\dagger}$ $71.5(\pm 2.4)^{\dagger}$ $85.4(\pm 1.4)^{\dagger}$
Not-E2E [85] ResNet-BN TCN, GRU 3 $92.0(\pm 5.3)$ $77.1(\pm 11.5)$
TMRNet [39] ResNeST LSTM, Attn. 2 $90.1(\pm 7.6)$ $79.1(\pm 5.7)$
PATG [40] SE-ResNet GNN 2 $91.4$
Zhang et al. [88] C3D LSTM 2 $85.9(\pm 7.9)$
He et al. [28] I3D GRU 2 $88.3(\pm 1.0)$
TeSTra [90] Transf. Transf. 2 $90.1$ $71.6$
He et al. [28] Video-Transf. GRU 2 $90.9(\pm 0.01)$
Dual Pyramid [12] Transf. Transf. 2 $91.4$ $75.4$
@@
LoViT [48] Transf. Transf. 2 $\it 92.4(\pm 6.3)$ $\it 81.2(\pm 9.1)$ $\it 91.5$ $\it 74.2$
SKiT [49] Transf. Transf. 2 $\it 93.4(\pm 5.2)$ $\it 82.6$ $\it 92.5$ $\it 76.7$
@@
Table 4: State of the art in surgical phase recognition on AutoLaparo [75] with 10/4/7 split. We report mean±$\pm$±std. over runs.
@@
SV-RCNet [37] $75.6$ $64.0$ $59.7$ $47.2$
TMRNet [39] $78.2$ $66.0$ $61.5$ $49.6$
TeCNO [13] $77.3$ $66.9$ $64.6$ $50.7$
Trans-SVNet [24] $78.3$ $64.2$ $62.1$ $50.7$
LoViT [48] $81.4$ $\bf 85.1$ $65.9$ $56.0$
SKiT [49] $82.9$ $81.8$ $70.1$ $59.9$
@@
Table 5: State of the art in surgical step recognition on CATARACTS [91] with a 25/25 split. Our models are trained on a 20/5/25 split to allow for model selection. State-of-the-art scores are printed bold and we underline scores which outperform all methods except Chen et al. [12].
@@
3D-CNN [21] $80.1$ $66.2$ $55.7$ $45.9$
SV-RCNet [37] $81.3$ $66.0$ $57.0$ $47.2$
TeCNO [13] $79.0$ $62.6$ $56.9$ $45.1$
Trans-SVNet [24] $77.8$ $61.3$ $55.0$ $43.8$
Dual Pyramid [12] $\bf 84.2$ $69.3$ $\bf 66.4$ $\bf 53.7$
@@
Cholec80: To be comparable to previous work [24, 37, 39], we retrain our models on a 32/8/40 split. As evaluation metrics, we report the mean video-wise accuracy and the mean video-wise balanced accuracy (i.e. mean video-wise macro recall) as well as variants of accuracy and Jaccard scores with relaxed boundaries, i.e. where errors at phase transitions are ignored under certain conditions. Relaxed metrics were proposed in the M2CAI 2016 Challenge [66, 70]. Table 3 shows our results.
Through partial freezing on 256-frame sequences, all our models with BN-free backbones (FrozenBN, GN and ConvNeXt) outperform the previous state of the art, including complex multi-stage methods with attention-based temporal models [24, 39, 12], complex augmentation pipelines [85] or Transformer backbones [12, 28]. Models trained completely end to end (E2E) on shorter 64-frame sequences perform slightly worse. However, the ConvNeXt variant still achieves SOTA performance and the GN variant beats or matches other ResNet-based methods. Note that Not-E2E [85] achieves slightly higher relaxed accuracy but much lower relaxed Jaccard. The completely end-to-end FrozenBN model underperforms, as already observed in Hypothesis 3.